# Using LLMs in Generating Design Rationale for Software Architecture Decisions

### v1

@@ 3.5. Evaluation Methods
We employed both accuracy metrics and IHUM-category to evaluate the generated DR. Accuracy metrics, including Precision, Recall, and F1-score, are employed for quantitative analysis. IHUM-category, a classification we defined in this study, is used for the qualitative analysis of DR, which is detailed in Section 3.5.2. The pilot evaluation was conducted by the first and fourth authors. They randomly selected 15 DR generated by LLMs to conduct a pilot evaluation: five generated with zero-shot, five with CoT, and five with LLM-based agents. The inter-rater reliability between the two authors was measured by the Cohen’s Kappa coefficient (Cohen, 1960). The Cohen’s Kappa value for the accuracy metrics is 0.863, whereas for IHUM-category is 0.738, indicating almost perfect and substantial agreement between the two authors, respectively. Then, the second and third authors joined the discussion with the first and fourth authors to resolve discrepancies and reach a consensus on the evaluation criteria. Subsequently, the first author conducted a formal evaluation on all generated DR. The evaluation results were thoroughly reviewed by the four authors (the first to fourth authors), with discrepancies resolved through discussion to reach a consensus. It is important to note that we did not use text similarity and semantic similarity metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and METEOR (Banerjee and Lavie, 2005), to evaluate the LLM-generated DR. Based on the pilot experiments conducted on 10 randomly selected DR generated by LLMs, we observed a significant difference between the text provided by human experts and that generated by LLMs. Simply applying text and semantic similarity metrics yields very low values with all the three prompting strategies, making it challenging to draw valid experimental conclusions.
@@ 3.5.1. Accuracy Metrics
For accuracy metrics, we used Precision, Recall, and F1-score to evaluate the alignment of DR generated by LLMs with those provided by human experts. The three metrics are commonly used to measure the quality of DR (e.g., (Zhao et al., 2024; Soliman and Keim, 2025)). As mentioned in Section 2, DR comprises a set of arguments. We compared the arguments provided in the DR generated by LLMs with those extracted from human experts during the data extraction phase. We then labeled the arguments into true positive (TP), false positive (FP), and false negative (FN). To be specific, we set the following criteria for labeling arguments:
@@
Equations  (2),  (3) and  (4) present three accuracy metrics, Precision, Recall, and F1-score. For a given data point $(P_{i},D_{i},DR_{i})$, $TP_{i}$, $FP_{i}$, and $FN_{i}$ are the number of arguments classified as true positive, false positive, and false negative in the DR. As the number of arguments from human experts is typically small (usually less than three), slight variations in the classification of arguments as $TP$, $FP$, and $FN$ can cause significant fluctuations in the metrics. For instance, when two arguments are provided by an expert, an increase in $TP_{i}$ from one to two results in a sharp rise in Recall from 50% to 100%. Therefore, these accuracy metrics are derived by calculating the weighted average based on the number of arguments provided by human experts for 100 data points.
(2) $\text{Precision}=\frac{\sum(TP_{i})}{\sum(TP_{i}+FP_{i})}$
(3) $\text{Recall}=\frac{\sum(TP_{i})}{\sum(TP_{i}+FN_{i})}$
[…]
@@ 3.5.2. IHUM-Category
We introduced the IHUM-category to qualitatively assess the argument points in DR generated by LLMs. Accuracy metrics are used to compare the similarity between the DR generated by LLMs and that provided by human experts. However, there may be a chance that the responses generated by LLMs do not align with those of human experts (i.e., FN); the arguments still offer valuable insights and provide alternative perspectives. We proposed that the evaluation of LLM-generated DR could follow a methodology akin to the one employed by Mahajan et al. (Mahajan et al., 2020) for assessing responses on SO. Therefore, we introduced the IHUM-category, which is adapted from the metrics in (Mahajan et al., 2020), with adjustments made to effectively assess DR. IHUM-category is used to manually classify all the argument points, allowing for a more comprehensive evaluation of the DR generated by LLMs. The definition of IHUM-category is presented in Table 3.
@@ 4. Results and Interpretation
In this section, we present the results of the three RQs with their interpretation. Specifically, we provide the accuracy metrics and IHUM-category classification results of the DR generated by the five LLMs (see Section 3.2) using three distinct prompting strategies (see Section 3.4), detailed in Tables 45, and 6, respectively. Table 7 presents the average performance of the five LLMs in generating DR with three prompting strategies. In Tables 4 to 7, the best-performing results among the five LLMs are highlighted using both underlining and bold text. Besides, since the number and proportion of arguments in the “Uncertain” and “Misleading” categories are negative indicators, the smallest values are highlighted in the tables.
@@ 4.1.1. For Accuracy Metrics
The average Precision, Recall, and F1-score of the DR generated by the five LLMs with the zero-shot method are 0.278, 0.627, and 0.381, respectively. The results show that approximately one-quarter of the arguments in the DR generated by the LLMs align with those provided by human experts. Additionally, over 60% of the expert-provided arguments are also mentioned in the LLM-generated DR. Among the five LLMs, gemini-1.0-pro exhibits the highest Precision (0.295) and F1-score (0.406), while gpt-4-0613 achieves the highest Recall (0.697). The relatively lower Precision and F1-score are mainly due to the higher number of arguments in the generated DR, which is approximately 2.3 times greater than those provided by human experts.
@@ 4.1.2. For IHUM-Category
The average proportion of arguments classified under “Insightful”, “Helpful”, “Uncertain”, and “Misleading” categories are 27.44%, 64.45%, 4.87%, and 3.24%, respectively. It is worth noting that besides the 27.44% arguments that are consistent with human experts (“Insightful”), around 65% are not considered by the human experts, but also offer “Helpful” insights into architectural design. Additionally, 4.87% of the arguments are indeterminate (“Uncertain”) and 3.24% contain potential errors (“Misleading”), representing a relatively small proportion. Among the five LLMs, gemini-1.0-pro generated the highest proportion of “Insightful” arguments (29.45%), while gpt-4-0613 produces the highest proportion of “Helpful” arguments (69.44%), along with the lowest proportions in the “Uncertain” (4.16%) and “Misleading” categories (1.45%).
@@
Precision 0.280 0.231 0.224 0.235 0.214 0.237
Recall 0.682 0.717 0.657 0.667 0.697 0.684
F1-score 0.396 0.350 0.335 0.347 0.327 0.351
Argument count 483 614 578 562 646 576.6
Insightful (I) 135 142 130 132 138 135.4
Helpful (H) 314 442 400 389 463 401.6
Uncertain (U) 25 23 30 23 30 26.2
Misleading (M) 9 7 18 18 15 13.4
I/(I+H+U+M) 27.95% 23.13% 22.49% 23.49% 21.36% 23.68%
H/(I+H+U+M) 65.01% 71.99% 69.20% 69.22% 71.67% 69.42%
U/(I+H+U+M) 5.18% 3.75% 5.19% 4.09% 4.64% 4.57%
M/(I+H+U+M) 1.86% 1.14% 3.11% 3.20% 3.32% 2.33%
@@ 4.2.1. For Accuracy Metrics
Except for gpt-3.5-turbo, the performance of the other four LLMs in terms of Precision and F1-score deteriorates with the CoT method, compared to the performance of the zero-shot method. This can be explained by the significantly higher number of arguments generated in the DR when using the CoT template, reaching 5.76 per DR on average, compared to 4.55 per DR with the zero-shot method. The increased number of arguments stems from the CoT template explicitly guiding LLMs to conduct a trade-off analysis after providing both the advantages and disadvantages. In contrast, the DR generated with the zero-shot template is less likely to consider the negative impacts of architecture decisions. Due to the increase in the number of arguments generated with the CoT template, the DR considers more diverse perspectives compared to the zero-shot method, which leads to a 9.1% increase in Recall. Additionally, the decrease in Precision and F1-score is also primarily attributed to the increase in the total number of arguments generated in the DR, which leads to a lower proportion of arguments consistent with those provided by human experts. However, this does not imply a decline in the quality of DR generated using the CoT method, as we cannot expect the LLM-generated DR to consist solely of arguments provided by human experts. Arguments not considered by human experts in the LLM-generated DR can also potentially offer valuable insights, which are further evaluated according to the IHUM category in Section 4.2.2.
@@ 4.2.2. For IHUM-Category
Table 7 presents a comparison of the average performance between the zero-shot and CoT methods. Although the proportion of “Insightful” arguments decreases, the absolute number of “Insightful” arguments generated with the CoT method increases, which is also reflected in the higher average Recall. Additionally, compared to the zero-shot method, the number of “Helpful” arguments increases by 36.3%, while the proportion increases by 7.7%. This improvement indicates that, with the CoT method, LLMs not only generate more “Insightful” arguments but also present a broader range of “Helpful” arguments, offering architects a more comprehensive perspective on architecture decisions. However, in certain DR, such as the one generated by gpt-4-0613 for SO #78536441 and mistral-7B for Discussion #9022, we identified an excess of arguments, which are 14 and 10 respectively. Some arguments that are less relevant to the architecture decision may result in the DR appearing overly lengthy and unfocused. Potential issues of the CoT method are further discussed in Section 5.1.
The proportion of “Uncertain” arguments has not significantly decreased. However, the proportion of “Misleading” arguments in the DR has declined. It is noticed that the CoT method enhances reasoning process by outlining the strengths and drawbacks of architecture decisions for conducting a trade-off analysis, thereby mitigating some of the errors observed in the zero-shot approach. For example, SO #77745340 examines the rationale for “running .NET projects targeting different framework versions in separate processes”, with Inter Process Communication (IPC) mechanisms implied. The DR generated by gpt-3.5-turbo using the zero-shot method suggests that “running components in separate processes can utilize system resources more efficiently”, concluding that the whole system performance is enhanced. However, the introduction of IPC actually increases the system’s performance costs due to the need for serialization and deserialization during data transmission between different processes. While the DR generated with the zero-shot method ignores this issue (the cost of IPC), the DR generated by the CoT method considers the issue when analyzing the disadvantages of this architecture, ultimately leading to a more accurate trade-off conclusion.
@@
Precision 0.291 0.255 0.269 0.254 0.266 0.267
Recall 0.717 0.747 0.707 0.687 0.717 0.715
F1-score 0.414 0.380 0.389 0.371 0.386 0.389
Argument count 488 580 521 535 533 531.4
Insightful (I) 142 148 140 136 142 141.6
Helpful (H) 318 407 345 365 363 359.6
Uncertain (U) 21 20 26 21 21 21.8
Misleading (M) 7 5 10 13 7 8.4
I/(I+H+U+M) 29.10% 25.52% 26.87% 25.42% 26.64% 26.71%
H/(I+H+U+M) 65.16% 70.17% 66.22% 68.22% 68.11% 67.58%
U/(I+H+U+M) 4.30% 3.45% 4.99% 3.93% 3.94% 4.12%
M/(I+H+U+M) 1.43% 0.86% 1.92% 2.43% 1.31% 1.59%

### v3

@@ 3.5. Evaluation Methods
According to Equation 1, each DR consists of multiple arguments $\{\bm{A}_{1},\bm{A}_{2},\ldots,\bm{A}_{n}\}$. Using individual arguments in DR as the fundamental unit of analysis, we applied both quantitative accuracy metrics and the qualitative IHUM-category to comprehensively evaluate the generated DR. Accuracy metrics, including Precision, Recall, and F1-score, are employed for quantitative analysis. IHUM-category, a classification we defined in this study, is used for the qualitative analysis of DR, which is detailed in Section 3.5.2. The pilot evaluation was conducted by the first and fourth authors. They randomly selected 15 DR generated by LLMs to conduct a pilot evaluation: five generated with zero-shot, five with CoT, and five with LLM-based agents. The inter-rater reliability between the two authors was measured by the Cohen’s Kappa coefficient (jacob1960coefficient). The Cohen’s Kappa value for the accuracy metrics is 0.863, whereas for IHUM-category is 0.738, indicating almost perfect and substantial agreement between the two authors, respectively. Then, the second and third authors joined the discussion with the first and fourth authors to resolve discrepancies and reach a consensus on the evaluation criteria. Subsequently, the first author conducted a formal evaluation of all generated DR. The evaluation results were thoroughly reviewed by the four authors (the first to fourth authors), with discrepancies resolved through discussion to reach a consensus. It is important to note that we did not use text similarity and semantic similarity metrics such as BLEU (papineni2002bleu), ROUGE (lin2004rouge), and METEOR (banerjee2005meteor), to evaluate the LLM-generated DR. Based on the pilot experiments conducted on 10 randomly selected DR generated by LLMs, we observed a significant difference between the text provided by human experts and that generated by LLMs. Simply applying text and semantic similarity metrics yields very low values with all three prompting strategies, making it challenging to draw valid experimental conclusions.
@@ 3.5.1. Accuracy Metrics
For accuracy metrics, we used Precision, Recall, and F1-score to evaluate the alignment of DR generated by LLMs with those provided by human experts. The three metrics are commonly used to measure the quality of DR (e.g., (zhao2024drminer; soliman2025do)). As mentioned in Section 2, DR comprises a set of arguments. We compared the arguments provided in the DR generated by LLMs with those extracted from human experts during the data extraction phase. We then labeled the arguments into true positive (TP), false positive (FP), and false negative (FN). To be specific, we set the following criteria for labeling arguments:
@@
Equations  (2),  (3) and  (4) present three accuracy metrics, Precision, Recall, and F1-score. For a given data point $(P_{i},D_{i},DR_{i})$, $TP_{i}$, $FP_{i}$, and $FN_{i}$ are the number of arguments classified as true positive, false positive, and false negative in the DR. As the number of arguments from human experts is typically small (usually less than three), slight variations in the classification of arguments as $TP$, FPFP, and $FN$ can cause significant fluctuations in the metrics. For instance, when two arguments are provided by an expert, an increase in $TP_{i}$ from one to two results in a sharp rise in Recall from 50% to 100%. Therefore, these accuracy metrics represent a weighted average over the 100 data points, with weights determined by the number of arguments provided in each DR. For example, weighted Precision is calculated by averaging the per-DR precision scores, weighted by the number of predicted arguments (i.e., $TP_{i}$$FP_{i}$) in each DR.
(2) $\text{Precision}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum\left(\frac{TP_{i}+FP_{i}}{\sum(TP_{j}+FP_{j})}\frac{TP_{i}}{TP_{i}+FP_{i}}\right)}=\frac{\sum(TP_{i})}{\sum(TP_{i}+FP_{i})}$
(3) $\text{Recall}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum\left(\frac{TP_{i}+FN_{i}}{\sum(TP_{j}+FN_{j})}\frac{TP_{i}}{TP_{i}+FN_{i}}\right)}=\frac{\sum(TP_{i})}{\sum(TP_{i}+FN_{i})}$
(4) $\text{F1-score}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
@@ 3.5.2. IHUM-Category
We introduced the IHUM-category to qualitatively assess the argument points in DR generated by LLMs. Accuracy metrics are used to compare the similarity between the DR generated by LLMs and that provided by human experts. However, there may be a chance that the responses generated by LLMs do not align with those of human experts (i.e., FN); the arguments still offer valuable insights and provide alternative perspectives. We proposed that the evaluation of LLM-generated DR could follow a methodology akin to the one employed by Mahajan et al. (mahajan2020recommending) for assessing responses on SO. Therefore, we introduced the IHUM-category, which is adapted from the metrics in (mahajan2020recommending), with adjustments made to effectively assess DR. IHUM-category is used to manually classify all the argument points, allowing for a more comprehensive evaluation of the DR generated by LLMs. The definition of IHUM-category is presented in Table 4.
@@
As part of our evaluation, we calculated the proportion of each of the four IHUM categories and used them as assessment metrics. Equation 5 illustrates the computation of the proportion of “Insightful” arguments, with analogous formulas applied to the other categories (“Helpful”, “Uncertain”, and “Misleading”). To maintain consistency with the accuracy metrics, a weighted calculation based on the number of arguments is employed.
(5) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{I}{I+H+U+M}=\sum\left(\frac{I_{i}+H_{i}+U_{i}+M_{i}}{\sum(I_{j}+H_{j}+U_{j}+M_{j})}\frac{I_{i}}{I_{i}+H_{i}+U_{i}+M_{i}}\right)=\frac{\sum(I_{i})}{\sum(I_{i}+H_{i}+U_{i}+M_{i})}$}
3.6. Interviews with Practitioners
To answer RQ4, we conducted semi-structured interviews to understand the practitioners’ perceptions of the LLM-generated DR by following the guidelines proposed by Wohlin et al. (Wohlin2012ESE). Semi-structured interviews enable practitioners to talk deeply and freely about their perspectives on the DR generated by LLMs based on their experience. The interview consists of three steps:
  • Step 1: Before the interviews, participants were introduced to the research objectives and methodology of our study. They were then asked to review five randomly selected samples from our dataset (dataset), comprising architecture problems and decisions, along with the corresponding DR provided by human experts and generated by LLMs. The five samples were compiled in a PDF file and distributed to the participants for review prior to the interviews. The PDF file that contains the five samples has been provided in our dataset (dataset).
  • Step 2: The interviews were conducted by the first author. Each participant provided a brief overview of their background, including their professional role, as well as their years of experience in software development and architecture design.
  • Step 3: During the semi-structured interviews, practitioners responded to five Interview Questions (IQs). With participants’ consent, the interviews were audio-recorded to facilitate subsequent analysis.
The five IQs were designed to explore two key aspects of LLM-generated DR from the perspective of practitioners: trustworthiness and applicability. Specifically, IQ1.1 and IQ1.2 focus on participants’ level of trust in the DR produced by LLMs and the types of information that could enhance their trust. IQ2.1 to IQ2.3 examine the potential use of LLM-generated DR in practice, including possible application scenarios and perceived challenges. The five IQs are as follows:
  • IQ1.1: To what extent would you trust the DR generated by LLMs for architecture decisions? Why? (Options range from “Do not trust at all” to “Fully trust without human validation”.)
  • IQ1.2: What types of information (e.g., project context, domain-specific knowledge) should be provided to LLMs to enhance the trustworthiness of the generated DR for architecture decisions?
  • IQ2.1: Would you consider using LLM-generated DR in practice? (Options range from “Definitely not” to “Definitely yes”.)
  • IQ2.2: What scenarios (e.g., comparing architectural alternatives, understanding architecture design) might be suitable for applying LLM-generated DR for architecture decisions?
  • IQ2.3: What are the main challenges or limitations in applying LLM-generated DR in practice?
@@ 4. Results and Interpretation
In this section, we present the results of the four RQs with their interpretation. Specifically, we provide the accuracy metrics and IHUM-category classification results of the DR generated by the five LLMs (see Section 3.2) using three distinct prompting strategies (see Section 3.4), detailed in Tables 56, and 7, respectively. Table 5 to Table 7 additionally report the average performance of the metrics for DR generation, highlighting how LLMs perform under each prompting strategy and supports the interpretation of results in Section 4 as well as the discussion in Section 5. Table 8 presents the average performance of the five LLMs in generating DR with three prompting strategies. In Table 5 to Table 8, the best-performing results of the five LLMs are highlighted using bold and underlined text. Notably, since the number and proportion of arguments in the “Uncertain” and “Misleading” categories are negative indicators, the smallest values are highlighted in the tables.
@@ 4.1. RQ1: How effective are LLMs in generating and recovering design rationale for architecture decisions?
Note: Best results among the five LLMs are highlighted using underlining and bold font, such as xx.xx%.
@@ 4.1.1. For Accuracy Metrics
The average Precision, Recall, and F1-score of the DR generated by the five LLMs with the zero-shot method are 0.278, 0.627, and 0.381, respectively. Among the five LLMs, gemini-1.0-pro exhibits the highest Precision (0.295) and F1-score (0.406), while gpt-4-0613 achieves the highest Recall (0.697).
The results of the zero-shot method show the capability of LLMs in generating DR using a basic prompting strategy. According to the accuracy results, about one in four arguments generated by the LLMs are consistent with those from human experts. Additionally, over 60% of the expert-provided arguments are also mentioned in the LLM-generated DR. The five LLMs exhibited relatively low precision, which in turn contributed to the low F1-score. However, this is primarily due to the tendency of LLMs to generate a larger number of arguments. In fact, the average number of arguments generated by the LLMs with zero-shot (4.55) is 2.3 times greater than that provided by human experts (1.98). Therefore, even though the LLMs’ generated DR covers most of the arguments provided by human experts, the Precision still does not exceed 0.3. Similar patterns are observed in both the CoT and LLM-based agent methods. The arguments generated by LLMs but not considered by human experts are further evaluated using the IHUM-category, as presented in Section 4.1.1. However, an issue of zero-shot method lies in that the arguments generated by LLMs tend to emphasize the advantages of the chosen architecture while neglecting its potential disadvantages. As a consequence, when human experts focus on architectural shortcomings in some architecture problems (e.g., Issue #33070 and Discussion #214), the Precision of the DR produced by LLMs under the zero-shot method is low. The limitation due to the absence of architectural shortcomings of the zero-shot method is further discussed in Section 5.1.
@@ 4.1.2. For IHUM-Category
The average proportion of arguments classified under “Insightful”, “Helpful”, “Uncertain”, and “Misleading” categories are 27.44%, 64.45%, 4.87%, and 3.24%, respectively. It is worth noting that besides the 27.44% arguments that are consistent with human experts (“Insightful”), around 65% are not considered by the human experts, but also offer “Helpful” insights into architecture design. Additionally, 4.87% of the arguments are indeterminate (“Uncertain”) and 3.24% contain potential errors (“Misleading”), representing a relatively small proportion. Among the five LLMs, gemini-1.0-pro generated the highest proportion of “Insightful” arguments (29.45%), while gpt-4-0613 produces the highest proportion of “Helpful” arguments (69.44%), along with the lowest proportions in the “Uncertain” (4.16%) and “Misleading” categories (1.45%).
Among the three prompting strategies, the DR generated by the zero-shot method exhibits the highest proportion of “Misleading” arguments. This result may be attributed to the prompt template of zero-shot, which directly instructs the LLMs to generate DR without incorporating an explicit reasoning process. The absence of such guided reasoning contributes to the lack of trade-off analysis in the DR. Besides, in addition to the arguments that can be categorized using the IHUM-category, the DR generated by the zero-shot method occasionally contains content that is unrelated to the rationale of architecture decisions (e.g., alternative architectures, code snippets). This tendency is especially evident in the DR generated by gpt-4-0613 and llama3-8B, which contain a higher proportion of such extraneous information than other LLMs. The limitations of zero-shot are further discussed in Section 5.1.
@@
gpt-3.5-turbo gpt-4-0613 gemini-1.0-pro llama3-8B mistral-7B Average zero-shot
Precision 0.280 0.231 0.224 0.235 0.214 0.237 0.278
Recall 0.682 0.717 0.657 0.667 0.697 0.684 0.627
F1-score 0.396 0.350 0.335 0.347 0.327 0.351 0.381
Argument count 483 614 578 562 646 576.6 455
Insightful (I) 135 142 130 132 138 135.4 124.2
Helpful (H) 314 442 400 389 463 401.6 294.6
Uncertain (U) 25 23 30 23 30 26.2 22
Misleading (M) 9 7 18 18 15 13.4 14.2
I/(I+H+U+M) 27.95% 23.13% 22.49% 23.49% 21.36% 23.68% 27.44%
H/(I+H+U+M) 65.01% 71.99% 69.20% 69.22% 71.67% 69.42% 64.45%
U/(I+H+U+M) 5.18% 3.75% 5.19% 4.09% 4.64% 4.57% 4.87%
M/(I+H+U+M) 1.86% 1.14% 3.11% 3.20% 3.32% 2.33% 3.24%
Note: Best performance among the five LLMs are highlighted using underlining and bold font, such as xx.xx%.
@@ 4.2.1. For Accuracy Metrics
Except for gpt-3.5-turbo, the performance of the other four LLMs in terms of Precision and F1-score deteriorates with the CoT method, compared to the performance of the zero-shot method. This can be explained by the higher number of arguments generated in the DR when using the CoT template, reaching 5.76 per DR on average, compared to 4.55 per DR with the zero-shot method. Among the five LLMs, mistral-7B exhibits the largest increase in the number of arguments (+2.0), followed by llama3-8B (+1.6) and gemini-1.0-Pro (+1.4), while gpt-4-0613 (+0.6) and gpt-3.5-turbo (+0.4) show relatively smaller increases. The increased number of arguments stems from the CoT template explicitly guiding LLMs to conduct a trade-off analysis after providing both the advantages and disadvantages. In contrast, the DR generated with the zero-shot template is less likely to consider the negative impacts of architecture decisions. Due to the increase in the number of arguments generated with the CoT template, the DR considers more diverse perspectives compared to the zero-shot method, which leads to a 9.1% increase in Recall. Additionally, the decrease in Precision and F1-score is also primarily attributed to the increase in the total number of arguments generated in the DR, which leads to a lower proportion of arguments consistent with those provided by human experts. However, this does not imply a decline in the quality of DR generated using the CoT method, as we cannot expect the LLM-generated DR to consist solely of arguments provided by human experts. Arguments not considered by human experts in the LLM-generated DR can also potentially offer valuable insights, which are further evaluated according to the IHUM-category in Section 4.2.2.
@@ 4.2.2. For IHUM-Category
Table 6 also presents a comparison of the average performance between the zero-shot and CoT methods. Although the proportion of “Insightful” arguments decreases, the absolute number of “Insightful” arguments generated with the CoT method increases, which is also reflected in the higher average Recall. Additionally, compared to the zero-shot method, the number of “Helpful” arguments increases by 36.3%, while the proportion increases by 7.7%. This improvement indicates that, with the CoT method, LLMs not only generate more “Insightful” arguments but also present a broader range of “Helpful” arguments, offering architects a more comprehensive perspective on architecture decisions. However, in certain DR, such as the one generated by gpt-4-0613 for SO #78536441 and mistral-7B for Discussion #9022, we identified an excess of arguments, which are 14 and 10 respectively. Some arguments that are less relevant to the architecture decision may result in the DR appearing overly lengthy and unfocused. Potential issues of the CoT method are further discussed in Section 5.1.
The proportion of “Uncertain” arguments has not decreased. However, the proportion of “Misleading” arguments in the DR has declined. It is noticed that the CoT method enhances the reasoning process by outlining the strengths and drawbacks of architecture decisions for conducting a trade-off analysis, thereby mitigating some of the errors observed in the zero-shot approach. For example, SO #77745340 examines the rationale for “running .NET projects targeting different framework versions in separate processes”, with Inter Process Communication (IPC) mechanisms implied. The DR generated by gpt-3.5-turbo using the zero-shot method suggests that “running components in separate processes can utilize system resources more efficiently”, concluding that the whole system performance is enhanced. However, the introduction of IPC actually increases the system’s performance costs due to the need for serialization and deserialization during data transmission between different processes. While the DR generated with the zero-shot method ignores this issue (the cost of IPC), the DR generated by the CoT method considers the issue when analyzing the disadvantages of this architecture, ultimately leading to a more accurate trade-off conclusion. Besides, with the CoT method, the reasoning steps are explicitly specified in the prompt template, leading all LLMs to adopt a step-by-step structure when generating DR. As a result, the DR produced by the five LLMs exhibits a high degree of consistency in content organization. Notably, gemini-1.0-pro, llama3-8B, and mistral-7B tend to include a concluding summary at the end of the DR.
@@
gpt-3.5-turbo gpt-4-0613 gemini-1.0-pro llama3-8B mistral-7B Average zero-shot
Precision 0.291 0.255 0.269 0.254 0.266 0.267 0.278
Recall 0.717 0.747 0.707 0.687 0.717 0.715 0.627
F1-score 0.414 0.380 0.389 0.371 0.386 0.389 0.381
Argument count 488 580 521 535 533 531.4 455
Insightful (I) 142 148 140 136 142 141.6 124.2
Helpful (H) 318 407 345 365 363 359.6 294.6
Uncertain (U) 21 20 26 21 21 21.8 22
Misleading (M) 7 5 10 13 7 8.4 14.2
I/(I+H+U+M) 29.10% 25.52% 26.87% 25.42% 26.64% 26.71% 27.44%
H/(I+H+U+M) 65.16% 70.17% 66.22% 68.22% 68.11% 67.58% 64.45%
U/(I+H+U+M) 4.30% 3.45% 4.99% 3.93% 3.94% 4.12% 4.87%
M/(I+H+U+M) 1.43% 0.86% 1.92% 2.43% 1.31% 1.59% 3.24%
Note: Best performance among the five LLMs are highlighted using underlining and bold font, such as xx.xx%.